OpenStudy (anonymous):

Melissa is 24 years younger than joyce. In 2 years, Joyce will be 3 times as old as Melissa will be then. How old are they now?

OpenStudy (anonymous):

m = j - 24 j = m +24 3(m + 2) = j + 2 // Setting up our formulas from given info 3m+6 = j +2 //simplify 3m +6 = j +2 3m + 6 = m+24 +2 //substituting 3m + 6 = m +26 3m = m +20 3m - m = 20 2m = 20 m =10 m = j - 24 //take one of the first formulas //substitute 10 for m 10 = j -24 34 = j j = 34 10 and 34 in 2 years 12 and 36